Physical Chemistry

if an electron in hydrogen atom has an energy of -78.4 kilo calorie per mole. the orbit in which the electron is present is​:

Saumy Katiyar
8 Years agoGrade

1 Answer

Eshan
8 Years ago
Dear student,

Energy of one mole of electrons=-78.4kcal=-78.4\times 10^3\times 4.18J=-328025.6J

Therefore, energy of one such electron ineVis-\dfrac{328025.6}{6.023\times 10^{23}\times 1.6\times 10^{-19}}eV=-3.4eV

Therefore, the electron is in second orbit.